OpenStudy (anonymous):

given f(x) = x+ t and g(x) = 11x - 5, what is f(x) + g(x) a. 2x b. 11x c. 12x d. 55x-25

OpenStudy (cwrw238):

x + t + 11x - 5

OpenStudy (cwrw238):

you sure you've copied the question correctly?

OpenStudy (anonymous):

whoops. its suppose to be given f(x) = x + 5 and g(x) = 11x - 5, what is f(x) +g(x)

OpenStudy (anonymous):

Do the same as @cwrw238 did above.. :)

OpenStudy (solomonzelman):

\(\Large\color{black}{ g(x)=x+5}\) \(\Large\color{black}{ f(x)=11x-5}\) when you add \(\Large\color{black}{ f(x)}\) and \(\Large\color{black}{ g(x)}\) you are just adding what each of the functions is equal to. \(\Large\color{black}{ f(x)+g(x)}\) is , \(\Large\color{black}{ (x+5)~+~(11x-5)}\)

OpenStudy (solomonzelman):

All you have to do is to add like terms.

OpenStudy (anonymous):

ok thank you so much :)

OpenStudy (solomonzelman):

tell us what you get please, so that we can verify that you are correct.